# Anticipating Future with Large Language Model for Simultaneous Machine Translation

### v1

@@ Abstract
Simultaneous machine translation (SMT) takes streaming input utterances and incrementally produces target text. Existing SMT methods only use the partial utterance that has already arrived at the input and the generated hypothesis. Motivated by human interpreters’ technique to forecast future words before hearing them, we propose Translation by Anticipating Future (TAF), a method to improve translation quality while retraining low latency. Its core idea is to use a large language model (LLM) to predict future source words and opportunistically translate without introducing too much risk. We evaluate our TAF and multiple baselines of SMT on four language directions. Experiments show that TAF achieves the best translation quality-latency trade-off and outperforms the baselines by up to 5 BLEU points at the same latency (three words).
@@ 2 Related Works
xThough rule-based policies are easy to implement, learned adaptive policies demonstrate better quality-latency trade-offs. Ma et al. (2020b) proposed monotonic multihead attention to model the policy. Miao et al. (2021) developed a generative framework with a latent variable to make decisions. Zhang and Feng (2022) quantified the information weight transported from source to target and made decisions based on the amount of information received. Zhang and Feng (2023) modeled the simultaneous translation process as a hidden Markov model and optimized the likelihood of target sequence over multiple steps. Guo et al. (2024) further bridged the gap between SMT models and offline MT models. Yu et al. (2024) is the state-of-the-art adaptive policy that resolves the insufficient exploration issue of prior works by individually optimizing each source-target state.
@@ 4.3 Evaluation Metrics
We use SimulEval (Ma et al., 2020a) to evaluate our method and other baselines. We evaluate the quality of translation by comparing the hypothesis with the ground-truth translation using BLEU and COMET Guerreiro et al. (2024) [^1] https://huggingface.co/Unbabel/XCOMET-XXL. The latency is evaluated with Length Adaptive Average Lagging (LAAL) (Papi et al., 2022). Note that we treat the word (En, De) or character (Zh, Ja) as the unit during evaluation instead of the BPE token, following the practice of recent IWSLT competitions Ahmad et al. (2024). The latency calculated with the word or character is more intelligible than that computed with the BPE token and also enables fair comparison of models with different tokenizations.
@@ 5.2 Generalizable to Existing Pretrained MT models and LLMs
MT Model Size TAF Overhead
Transformer-Big 0.2B 8.07×8.07$\times$8.07 ×
MADLAD-400 3B 3.39×3.39$\times$3.39 ×
TowerInstruct 7B 1.04×1.04$\times$1.04 ×
Table 1: The additional computation overhead introduced by TAF on different MT backbones. LLM is fixed to 7B. As the size of the MT backbone increases, the additional overhead introduced by LLM prediction decreases.
@@ 5.3 Computation Cost
TAF introduces additional computation overhead to conventional SMT. We report the average wall-clock time of generating a full hypothesis with RALCP+TAF versus the base RALCP policy on different combinations of MT models and a 7B LLM. We run experiments on an A6000 GPU and an AMD EPYC 9354 32-Core CPU. Results are shown in Table 1. When using a small MT model Transformer-Big (0.2B), 7B LLM introduces 8.07×8.07$\times$8.07 × computational overhead. As the size of the MT model gets larger, the relative overhead reduces to 3.39×3.39$\times$3.39 × and 1.04×1.04$\times$1.04 ×.
@@ 5.4 Sensitivity to Hyperparameters
We examine how TAF is sensitive to the choice of hyperparameters. We evaluate on De-En direction and use RALCP+TAF. We vary the number of tokens to predict $l$, the number of continuations $n$, and the $k$ for top-$k$ sampling from $[2,4,6,8,10]$. We sweep $\tau$ from $0.5$ to $0.9$ with step size $0.1$ for each configuration.

### v2

@@ Abstract
Simultaneous machine translation (SMT) takes streaming input utterances and incrementally produces target text. Existing SMT methods mainly use the partial utterance that has already arrived at the input and the generated hypothesis. Motivated by human interpreters’ technique to forecast future words before hearing them, we propose Translation by Anticipating Future (TAF), a method to improve translation quality while retaining low latency. Its core idea is to use a large language model (LLM) to predict future source words and opportunistically translate without introducing too much risk. We evaluate our TAF and multiple baselines of SMT on four language directions. Experiments show that TAF achieves the best translation quality-latency trade-off and outperforms the baselines by up to 5 BLEU points at the same latency (three words) [^1] https://github.com/owaski/TAF.
@@ 2 Related Works
Though rule-based policies are easy to implement, learned adaptive policies demonstrate better quality-latency trade-offs. Ma et al. (2020b) proposed monotonic multihead attention to model the policy. Miao et al. (2021) developed a generative framework with a latent variable to make decisions. Zhang and Feng (2022) quantified the information weight transported from source to target and made decisions based on the amount of information received. Zhang and Feng (2023) modeled the simultaneous translation process as a hidden Markov model and optimized the likelihood of target sequence over multiple steps. Guo et al. (2024) further bridged the gap between SMT models and offline MT models. Yu et al. (2024) is the state-of-the-art adaptive policy that resolves the insufficient exploration issue of prior works by individually optimizing each source-target state.
@@ 4.3 Evaluation Metrics
We use SimulEval (Ma et al., 2020a) to evaluate our method and other baselines. We evaluate the quality of translation by comparing the hypothesis with the ground-truth translation using BLEU Papineni et al. (2002) and COMET Guerreiro et al. (2024) [^2] https://huggingface.co/Unbabel/XCOMET-XXL. The latency is evaluated with Length Adaptive Average Lagging (LAAL) (Papi et al., 2022). Note that we treat the word (En, De) or character (Zh, Ja) as the unit during evaluation instead of the BPE token, following the practice of recent IWSLT competitions Ahmad et al. (2024). The latency calculated with the word or character is more intelligible than that computed with the BPE token and also enables fair comparison of models with different tokenizations.
@@ 5.2 Generalizable to Existing Pretrained MT models and LLMs
We also demonstrate the perplexity of LLMs on the source text in Table 3 together with their latency and quality scores. The effectiveness of TAF is closely correlated with the LLM perplexity on the source text.
MT Model Size w/o TAF w/ TAF
Transformer-Big 0.2B 72ms 1067ms
MADLAD-400 3B 383ms 1339ms
TowerInstruct 7B 905ms 1506ms
Table 1: The average wall-clock time per step with and without TAF on different MT backbones. LLM is fixed to 7B. As the size of the MT backbone increases, the relative additional overhead introduced by LLM prediction decreases.
@@ 5.3 Computation Cost
TAF introduces additional computation overhead to conventional SMT. We report the average wall-clock time of generating a full hypothesis with RALCP+TAF versus the base RALCP policy on different combinations of MT models and a 7B LLM. We run experiments on an A6000 GPU and an AMD EPYC 9354 32-Core CPU. The results are shown in Table 1. When using a small MT model Transformer-Big (0.2B), 7B LLM introduces significant computational overhead. As the size of the MT model gets larger, the relative overhead reduces.
@@ 5.4 Sensitivity to Hyperparameters
We examine how TAF is sensitive to the choice of hyperparameters. We evaluate on De-En direction and use RALCP+TAF. We vary the number of tokens to predict $l$, the number of continuations $n$, and the $k$ for top-$k$ sampling from $[2,4,6,8,10]$. We sweep $\tau$ from $0.5$ to $0.9$ with step size $0.1$ for each configuration. Note that for $n=2$ and $\tau<1$, there is only one data point since RALCP always follows the first of $n=2$ translations.
@@ Appendix A Training Details of Offline MT
Appendix B LLM Perplexity on Source Text
To find out the correlation between LLM’s next token prediction and the final simultaneous translation performance, we measure the perplexity of MicroLlama-300m, Llama2-7b and Mistral-7b on source sentences of WMT20 En-Zh test set. We also measure the perplexity of Llama2-7b on unsegmented source documents as in Section 7. The results are shown in Table 3.